Suppose a condo generates $19,000 in cash flow in the first year. If the cash flows grow at 2% per year, the interest rate is 8%, and the building will be torn down in 20 years (the building is worthless after 20 years), what is the present value of the condos cash flow? Enter your response below (rounded to 2 decimal places). Number

Please show work!
Answer #1

Please find below the solution..

i ii iii iv=ii*iii
year Cash flow PVIF @ 8% present value
1    19,000.00               0.9259      17,592.59
2    19,380.00 =19000*102%               0.8573      16,615.23
3    19,767.60 =19380*102%               0.7938      15,692.16
4    20,162.95 =19767.60*102%               0.7350      14,820.37
5    20,566.21               0.6806      13,997.02
6    20,977.54               0.6302      13,219.41
7    21,397.09               0.5835      12,484.99
8    21,825.03               0.5403      11,791.38
9    22,261.53               0.5002      11,136.31
10    22,706.76               0.4632      10,517.62
11    23,160.89               0.4289        9,933.31
12    23,624.11               0.3971        9,381.46
13    24,096.59               0.3677        8,860.27
14    24,578.53               0.3405        8,368.03
15    25,070.10               0.3152        7,903.14
16    25,571.50               0.2919        7,464.08
17    26,082.93               0.2703        7,049.41
18    26,604.59               0.2502        6,657.77
19    27,136.68               0.2317        6,287.90
20    27,679.41               0.2145        5,938.57
present value    215,711.01
Therefore present value =       215,711.01